LASER SPECTROSCOPY OF EUROPIUM MONOFLUORIDE

\underline{MICHAEL J. DICK}, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, Ontario; PETER F. BERNATH, JIN-GUO WANG, JIAN TANG, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, Ontario.

A preliminary investigation into the spectroscopic properties of europium monofluoride has commenced. At present, high resolution spectra have been obtained for a series of transitions in the 570 - 670 nm region. Preliminary assignments indicate that these electronic transitions are originating in one of two low-lying electronic states spaced by approximately 1400 cm^{−1}. Comparison with previous work indicates that these states, assigned as the ground and first excited states, are both arising from same electron configuration of the europium ion (Eu^{+}(f^{7}s)). Results and analysis of the ongoing investigation will be presented and discussed in terms of the electron configurations of the various observed states.